OpenStudy (anonymous):

The theoretical probability of undesirable side effects resulting from taking Grebex is 2 in 13. If 260 people take Grebex to lower their blood pressure, how many will encounter undesirable side effects?

OpenStudy (anonymous):

\[\frac{2}{13}\times 260\]

OpenStudy (anonymous):

What do u think JakeV8???

OpenStudy (anonymous):

I think you are making this too hard... If I tell you that 2 out of 100 people get sick after taking a medicine, and 500 people took it, you would expect 2 out of each of 100 to get sick.... so there are 5 x 100 people, and 5 x 2 = 10 might get sick. So you multiply the total number of people by the probability of getting sick. 500 x (2/100) = 1000/100 = 10 Your problem has a probability of 2/13, and a population of 260. So you have: 260 x (2/13) you can reduce first, or multiply then reduce... your choice.

OpenStudy (anonymous):

Im lost could you walk me through my problem...

OpenStudy (anonymous):

260x(2x13)= 6760

OpenStudy (anonymous):

no, it should be 2/13, not 2x13 2 out of every 13 people will get sick. So if you have 13, 2 will get sick. If you have 26, that's twice 13, so twice as many will get sick... 4. If you have 130 people, 2 out of every 13 will get sick... there are ten groups of 13 in 130, so 2 people in each of 10 groups will get sick.... 20 sick people total. For 260 people, there are 20 groups of 13, since 260/13 = 20... so for each of the 20 groups, 2 of those people will get sick... 2 x 20 = 40 sick people.

OpenStudy (anonymous):

\[260 \times \frac{ 2 }{ 13 } = \frac{ 260 \times 2 }{13} = (\frac{ 260 }{13})(2) = (20)(2) = 40\]